Intro Finance ch 4

In 1896, the first Green Jacket Golf Championship was held. The winner’s prize money was $195. In 2018, the winner’s check was $2,430,000.


a.What was the annual percentage increase in the winner’s check over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
b.

If the winner’s prize increases at the same rate, what will it be in 2049? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 1,234,567.89.)
Answer #1

Part a:

The annual percentage increase = 8.04%

Part b:
Calculation for future value in 2049:
Number of years=2049-2018=31
Present value in year 2018=$2430000
Future value=Present value*(1+annual rate)^Number of years
=2430000*(1+8.04%)^31
=$26713335.25